Find two consecutive odd integers whose product is 195.
Found 2 solutions by Targetweek, sommersd:
Answer by Targetweek(62)

well 13*13 = 169 and 15*15 = 225
so i tried 13*15 = 195
your answer is 13 and 15

Answer by sommersd(1)

Deandre can choose to use Plan A or Plan B for his long distance charges. For each plan, cost (in dollars) is a function of minutes used (per month) as shown below.